I have tried to solve these type problems and I cannot figure out the steps to solve them, if someone could help me with this one and show me steps on how to solve others like it I would really appreciate it.
Tyler has two savings accounts that his grandparents opened for him. The two accounts pay 10% and 16% in annual interest; there is $500 more in the account that pays 16% than there is in the other account. If the total interest for a year is $288, how much money does he have in each account?
How much was invested at 10%_____ and 16%______
Thank you so very much
Sabrina

Answer by ankor@dixie-net.com(12692)

Tyler has two savings accounts that his grandparents opened for him. The two accounts pay 10% and 16% in annual interest; there is $500 more in the account that pays 16% than there is in the other account. If the total interest for a year is $288, how much money does he have in each account?
:
Let x = amt at 10%
:
The problems states that there is $500 more in the 16% account, therefore:
(x+500) = amt at 16%
:
The problem also states:
10% acc interest + 16% acc interest = $288
.10x + .16(x+500) = 288
:
Multiply what's inside the brackets:
.10x + .16x + 80 = 288
.26x = 288 - 80; subtracted 80 from both sides, added the x's
.26x = 208
x = 208/.26
x = $800 invested at 10%
:
800 + 500 = $1300 invested at 16%
:
:
Check our solutions:
.10(800) + .16(1300) =
80 + 208 = 288
:
How about this? Can you do these kind now?